OpenStudy (abbycross167):

Can someone please help me with this question? I will fan and give a medal... . If x = 2 and y = -3, what is the value of 2x^2 - 3xy – 2y^2? a. -20 b. -2 c. 8 d. 16

OpenStudy (abbycross167):

I have tried to answer it but I keep getting the wrong answer

OpenStudy (xapproachesinfinity):

can you show how you did it

OpenStudy (abbycross167):

2*2^2-3*2*-3-2*-3^2= 44

OpenStudy (astrophysics):

\[2(2)^2 -3(2)(-3)-2(-3)^2\] be careful, put brackets.

OpenStudy (astrophysics):

\[(-3)^2 = 9~~~~~ and~~~-(3)^2 = -9\] notice the difference

OpenStudy (anonymous):

2(2)^2-3(2)(-3)-2(-3)^2 2(4)-3(2)(-3)-2(9) 8+18-18 8

OpenStudy (abbycross167):

ooh ok! thank y'all vary much!!!